Find fx(0,0) and fx (x,y) ,where (x, y) ≠(0,0) for the following function

f (x,y)={xy^3/x^2+y^2 , (x,y)not=to (0,0)
0 (x,y)= to (0,0). }

Is fx continuous at (0,0) ? Justify your answer.
1
Expert's answer
2020-04-01T11:29:58-0400
"f (x,y)=\begin{cases}
 \frac{xy^3}{x^2+y^2} &\text{if } (x,y)\neq (0,0) \\
 0 &\text{if } (x,y)=(0,0)
\end{cases}"


Find "f_x(0,0)."


"f_x(0,0) = \lim\limits_{h\to0}\frac{f(h,0) - f(0,0)}{h} = \lim\limits_{h\to0}\frac{0}{h}=0"

Find "f_x(x,y) \text{ where } (x,y)\neq(0,0)."


"f_x(x,y) = \frac{y^3(x^2+y^2) - 2x^2y^3}{(x^2+y^2)^2} = \frac{(y^2-x^2)y^3}{(x^2+y^2)^2}"

Therefore, the partial derivative "f" with respest to "x" is



"f_x (x,y)=\begin{cases}
 \frac{(y^2-x^2)y^3}{(x^2+y^2)^2}&\text{if } (x,y)\neq (0,0) \\
 0 &\text{if } (x,y)=(0,0)
\end{cases}"

To check continuity, find the next limit.


"\lim\limits_{(x,y)\to(0,0)} f_x(x,y)"

Proceed to polar coordinates "x = r\cos\varphi, y=r\sin\varphi."



"\lim\limits_{(x,y)\to(0,0)} \frac{(y^2-x^2)y^3}{(x^2+y^2)^2} = \lim\limits_{r\to0} \frac{(r^2\sin^2\varphi - r^2\cos^2\varphi)r^3\sin^3\varphi}{(r^2\cos^2\varphi+r^2\sin^2\varphi)^2} \\
= \lim\limits_{r\to0} \frac{(\sin^2\varphi - \cos^2\varphi)r^5\sin^3\varphi}{r^4} = 0"

Therefore, "\lim\limits_{(x,y)\to(0,0)} f_x(x,y) = f(0,0)" and it means that "f_x" is continuous.